### v1

@@ Appendix B Observational effects and cosmological information gain
Previous works [e.g. 36, 35, 111, 112, 85] have sought to separate instrumental effects, such as the beam, from the pure signal, which in principle will provide a cleaner isolation of the cosmological information. This can be achieved with a simple deconvolution of the beam at the power spectrum level, so that Equation 32 becomes
@@
However, treating these effects as part of the covariance (i.e. deconvolving the beam from the measured signal and including its inverse in the noise) is not equivalent to including them directly in the signal model of the power spectrum multipoles, which are the realistic measurable from a 21cm intensity mapping survey. Appendix B in Bernal et al. [28] demonstrated that when the damping from the beam (or other instrumental effects) is placed entirely in the covariance, the variance of the power spectrum involves terms of the form $\int d\mu\,1/\mathcal{B}^{2}(k,\mu)$, which diverge as  $\mathcal{B}\,{\rightarrow}\,0$, causing covariance to asymptote to infinity for the multipoles. However, earlier works [126, 147] provide an important complementary warning: damping or smoothing terms, whether from redshift errors, instrumental beams, or non-linear suppressions, represent losses of information but can appear to break degeneracies and tighten constraints, if not carefully treated.
@@
In future work, we will explore this further with realistic simulations to validate that any recovered cosmological information is unbiased. As the MeerKAT beam is now being characterised with increasingly high precision [59], simulations will also provide the ideal environment to test how accurately the beam must be known and whether dedicated nuisance parameters are required to capture its residual uncertainty. This will ultimately frame the challenge of breaking the growth-rate–beam degeneracy, where we will explore the additional information available in, for example, clustering $\mu$-wedges [81, 54] or a tomographic analysis; leveraging the differing redshift evolution of the beam and the growth rate.

### v2

@@ Appendix B Observational effects and cosmological information gain
Previous works [e.g. Bull:2014rha, Bull:2015lja, Pourtsidou:2015mia, Pourtsidou:2016dzn, Karagiannis:2022ylq] have sought to separate instrumental effects, such as the beam, from the pure signal, which in principle will provide a cleaner isolation of the cosmological information. This can be achieved with a simple deconvolution of the beam at the power spectrum level, so that Equation 33 becomes
@@
However, treating these effects as part of the covariance (i.e. deconvolving the beam from the measured signal and including its inverse in the noise) is not equivalent to including them directly in the signal model of the power spectrum multipoles, which are the realistic measurable from a 21cm intensity mapping survey. Appendix B in Bernal:2019jdo demonstrated that when the damping from the beam (or other instrumental effects) is placed entirely in the covariance, the variance of the power spectrum involves terms of the form $\int d\mu\,1/\mathcal{B}^{2}(k,\mu)$, which diverge as  $\mathcal{B}\,{\rightarrow}\,0$, causing covariance to asymptote to infinity for the multipoles. However, earlier works [Seo:2007ns, Wang:2012bx] provide an important complementary warning: damping or smoothing terms, whether from redshift errors, instrumental beams, or non-linear suppressions, represent losses of information but can appear to break degeneracies and tighten constraints, if not carefully treated.
@@
In future work, we will explore this further with realistic simulations to validate that any recovered cosmological information is unbiased. As the MeerKAT beam is now being characterised with increasingly high precision [2023AJ....165...78D], simulations will also provide the ideal environment to test how accurately the beam must be known and whether dedicated nuisance parameters are required to capture its residual uncertainty. This will ultimately frame the challenge of breaking the growth-rate–beam degeneracy, where we will explore the additional information available in, for example, clustering $\mu$-wedges [Jennings:2015lea, Cunnington:2020mnn] or a tomographic analysis; leveraging the differing redshift evolution of the beam and the growth rate.